# references.bib

### v1

@@ Abstract
This paper studies the linear system identification problem in the general case where the disturbance is sub-Gaussian, correlated, and possibly adversarial. First, we consider the case with noncentral (nonzero-mean) disturbances for which the ordinary least-squares (OLS) method fails to correctly identify the system. We prove that the $l_{1}$-norm estimator accurately identifies the system under the condition that each disturbance has equal probabilities of being positive or negative. This condition restricts the sign of each disturbance but allows its magnitude to be arbitrary. Second, we consider the case where each disturbance is adversarial with the model that the attack times happen occasionally but the distributions of the attack values are completely arbitrary. We show that when the probability of having an attack at a given time is less than 0.5, the $l_{1}$-norm estimator prevails against any adversarial noncentral disturbances and the exact recovery is achieved within a finite time. These results pave the way to effectively defend against arbitrarily large noncentral attacks in safety-critical systems.
@@ I Introduction
The challenge of this problem is that the states are correlated with each other, even when the disturbances are assumed to be independent and identically distributed (i.i.d.). Early papers have shown that the ordinary least-squares (OLS) estimator $\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{2}^{2}$ overcomes this correlation and achieve the consistency and the convergence towards the true matrix [ljung1976ols, moore1978strong, lai1982least]. Recently, instead of the asymptotic properties, several studies have focused on non-asymptotic analyses and provided an error rate of $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ between the true matrix and OLS estimates under i.i.d. zero-mean Gaussian and sub-Gaussian disturbances after finite samples are collected [simchowitz2018learningwithout, jedra2020finite, sarkar2021finite]. These results were extended to use the OLS estimates for the linear quadratic control with unknown matrices, in a way that once a small enough estimation error is attained, the estimates are treated as the true matrices, achieving $O(\sqrt{T})$ regret [mania2019certainty, simchowitz2020naive]. While OLS is indeed a statistically optimal estimator in such i.i.d. cases, little is known about the system identification when the disturbances are correlated or possibly selected adversarially based on the past information.
@@
Recently, in the machine learning literature, online control problem with any bounded adversarial disturbance, called online nonstochastic control, has gained considerable attention, among which [hazan2020nonstochastic, simchowitz2020improper, chen2021blackbox] leveraged OLS method to recover the matrix and attained $O(T^{2/3})$ regret. However, this polynomial regret implies a substantial cumulative loss, leading to an impractical algorithm for real-world applications. To address more applicable situations, [feng2021adv, yalcin2023exact, zhang2024exact] assumed that adversarial attacks occur with probability less than 1, and exactly recover the true matrix with a non-smooth $l_{2}$-norm estimator defined by $\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{2}$. In this case, practicality is achieved since the regret does not grow after the exact recovery. However, they assumed that each nonzero adversarial attack follows a zero-mean sub-Gaussian distribution, which restricts the range of possible attacks.
@@
In this paper, we allow noncentral (nonzero-mean) sub-Gaussian adversarial disturbances and use the $l_{1}$-norm estimator $\operatorname*{arg\,min}_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{1}$ to identify the unknown matrix within a finite time. Our work presents the first result in the literature for the exact retrieval of the true matrix even in the presence of adversarial noncentral disturbances. Under the assumption that the probability of having an attack at a given time is $p<1$, we construct two scenarios of the adversarial disturbance structure:
@@
    symmetric disturbance around zero scaled by an amplifying random vector (see Assumption 5) with $p<1$,
@@
    any adversarial noncentral disturbance with $p<0.5$.
The first scenario is applicable to the domain where the positive and negative disturbances occur with equal chance, each of which can be scaled arbitrarily. For example, beliefs based on behavioral bias may cause fluctuations in the stock market [lo2004adaptive] or energy demand [werth2021energy], which can be explained by disturbances of arbitrary magnitude. A more practical situation arises in the second scenario: it applies when an extremely large attack with a nonzero mean occasionally affects the system, such as natural disaster on power systems [buldyrev2010cat, wang2016power, yan2020power, trakas2022st], unanticipated malicious cyber attacks [pasqualetti2013cyber, duo2022cyber], and others. In this paper, we will show that the $l_{1}$-norm estimator is indeed a robust estimator capable of successfully prevailing over such disturbances. Note that while the paper focuses on autonomous linear systems, it can be readily generalized to linear systems with inputs as well as nonlinear systems linearly parameterized by basis functions. This generalization can be achieved by adopting the formulation and technique studied in [zhang2024exact], but this will be left as future work due to space restrictions.
@@
Notation. Let $\mathbb{R}^{d}$ denote the set of $n$-dimensional vector and $\mathbb{R}^{d\times d}$ denote the set of $d\times d$ matrices. For a matrix $A$, $\|A\|_{2}$ denotes the operator norm of the matrix. For a vector $x$, $\|x\|_{1}$ denotes the $l_{1}$-norm of the vector and $\|x\|_{1}$ denotes the $l_{2}$-norm of the vector. $x^{T}$ denotes the tranpose of the vector $x$. $x>0$ means that all entries of $x$ are positive. For two vectors $x$ and $y$ of the same dimensions, let $x\circ y$ denote entry-wise product of the same dimensions as $x$ and $y$. Let $\mathbb{E}$ denote the expectation operator. For the event $\mathcal{E}$, let $\mathbb{P}(\mathcal{E})$ denote the probability of the event. Also, let $\mathbb{I}(\mathcal{E})$ denote the indicator of the event, meaning that $\mathbb{I}(\mathcal{E})=1$ if the event occurs and $\mathbb{I}(\mathcal{E})=0$ otherwise. We use $O(\cdot)$ for the big-O notation.
@@ II Preliminaries
Definition 1 (sub-Gaussian scalar variables)
@@
Note that the $\psi_{2}$-norm satisfies all properties of norm, including positive definiteness, homogeneity, and triangle inequality. We introduce the equivalent definition of sub-Gaussian variables in the case where $\mathbb{E}[w]=0$.
Definition 2 (sub-Gaussian scalar variables (MGF))
@@
$\mathbb{E}[\exp(\lambda w)]\leq\exp(\lambda^{2}\eta_{w}^{2})$ (3)
for all $\lambda\in\mathbb{R}$. For any sub-Gaussian variable that satisfies (1) and $\mathbb{E}[w]=0$, (3) is satisfied with $\eta_{w}\leq 6\sqrt{e}\|w\|_{\psi_{2}}$ [vershynin2018high].
To leverage the equivalent definition (3) only holding for zero-mean random variables, the centering lemma in [vershynin2018high] is useful. We specify the previously omitted absolute constant in the lemma below.
Lemma 1 (Centering lemma, [vershynin2018high])
If $w$ is a sub-Gaussian variable satisfying (1), then $w-\mathbb{E}[w]$ is also sub-Gaussian and it holds that
$\|w-\mathbb{E}[w]\|_{\psi_{2}}\leq\left(1+\sqrt{\frac{\pi}{\log 2}}\right)\|w\|_{\psi_{2}}.$ (4)
Real-world systems have a limit on their actuators so they cannot accept arbitrarily large inputs. This enables us to assume that each disturbance follows sub-Gaussian variables of which the tail event rarely occurs. Hoeffding’s inequality provides sharp bounds on the tail event.
Lemma 2 (Hoeffding’s inequality, [vershynin2018high])
Suppose that a random variable $w$ satisfies (3) with the parameter $\eta_{w}$. Then, for all $s>0$, we have
$\mathbb{P}(w\geq s)\leq\exp\Bigr{(}-\frac{s^{2}}{(2\eta_{w})^{2}}\Bigr{)},$ (5)
$\mathbb{P}(w\leq-s)\leq\exp\Bigr{(}-\frac{s^{2}}{(2\eta_{w})^{2}}\Bigr{)}.$ (6)
@@
Definition 3 (sub-Gaussian vector variables)
@@
Now, we provide the subdifferential of the $l_{1}$-norm used for the analysis of the $l_{1}$-norm estimator.
Definition 4 (Subdifferential of the $l_{1}$-norm)
Given a vector $x\in\mathbb{R}^{d}$ and its entries $x^{1},\dots,x^{d}$, the subdifferential of $\|x\|_{1}$ is denoted by $\partial\|x\|_{1}$ and defined as $[\text{sgn}(x^{1}),\dots,\text{sgn}(x^{d})]^{T}$ where
$\text{sgn}(z)=\begin{dcases*}1,&if $z>0$,\\ [-1,1],&if $z=0$,\\ -1,&if $z<0$.\end{dcases*}$ (8)
The decomposability of the subdifferential along the coordinates simplifies the analysis.
@@ III Problem Formulation
where $A^{*}\in\mathbb{R}^{d\times d}$ is the unknown true matrix, $x_{t}\in\mathbb{R}^{d}$ is the state, and $w_{t}\in\mathbb{R}^{d}$ is a disturbance injected into the system at time $t$. Our goal is to identify the true matrix $A^{*}$, given the state trajectory $x_{0},\dots,x_{T}$. We assume that $A^{*}$ has bounded operator norm and each $x_{0},w_{0},\dots,w_{T-1}$ is sub-Gaussian to prevent an unbounded growth of the system. However, we allow adversarial noncentral disturbances, meaning that an adversary can design the disturbance $w_{t}$ based on the previous information $\mathcal{F}_{t}=\{x_{0},w_{0},\dots,w_{t-1}\}$, and $\mathbb{E}[w_{t}~{}|~{}\mathcal{F}_{t}]$ is not necessarily zero. We formally present the relevant assumptions below.
Assumption 1 (Operator norm)
@@
Assumption 2 (Maximum sub-Gaussian norm)
There exists $\sigma_{w}>0$ such that $\|x_{0}\|_{\psi_{2}}\leq\sigma_{w}$ and $\|w_{t}\|_{\psi_{2}}\leq\sigma_{w}$ conditioned on $\mathcal{F}_{t}$ for all $t\geq 0$ and for all $\mathcal{F}_{t}$.
Our goal is to achieve a finite-time exact recovery, meaning that the error between the true matrix and the estimate should be exactly zero after a finite time. However, it is well-established that the lower bound of the error is $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ if disturbances follow i.i.d. zero-mean Gaussian distribution [simchowitz2018learningwithout]. Our disturbance structure includes this special distribution, and therefore the lower bound $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ does not allow an exact recovery without a further assumption also adopted by [feng2021adv, yalcin2023exact, zhang2024exact].
Assumption 3 (Probabilistic Attack)
Let $w_{t}^{i}$ denote the $i^{\text{th}}$ entry of the disturbance $w_{t}$. Each coordinate of the disturbance has a positive probability of being zero conditioned on $\mathcal{F}_{t}$, meaning that there exists $p<1$ such that
$\mathbb{P}(w_{t}^{i}=0~{}|~{}\mathcal{F}_{t})\geq 1-p$ (10)
holds for all $i=1,\dots,d$, as well as $t$ and $\mathcal{F}_{t}$.
Note that Assumption 3 implies that the system is not under attack at time $t$ with probability at least $1-p$. Now, we consider the following $l_{1}$-norm estimator at time $T$:
$\min_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{1},$ (11)
given a state trajectory $x_{0},\dots,x_{T}$. We will show that $l_{1}$-estimator successfully overcomes noncentral distributions and achieves a finite-time exact recovery.

### v5

@@ Abstract
This paper studies the linear system identification problem in the general case where the disturbance is sub-Gaussian, correlated, and possibly adversarial. First, we consider the case with noncentral (nonzero-mean) disturbances for which the ordinary least-squares (OLS) method fails to correctly identify the system. We prove that the $l_{1}$-norm estimator accurately identifies the system under the condition that each disturbance has equal probabilities of being positive or negative. This condition restricts the sign of each disturbance but allows its magnitude to be arbitrary. Second, we consider the case where each disturbance is adversarial with the model that the attack times happen occasionally but the distributions of the attack values are arbitrary. We show that when the probability of having an attack at a given time is less than 0.5 and each attack spans the entire space in expectation, the $l_{1}$-norm estimator prevails against any adversarial noncentral disturbances and the exact recovery is achieved within a finite time. These results pave the way to effectively defend against arbitrarily large noncentral attacks in safety-critical systems.
@@ I Introduction
The challenge of this problem is that the states are correlated with each other, even when the disturbances are assumed to be independent and identically distributed (i.i.d.). Early papers have shown that the ordinary least-squares (OLS) estimator $\operatorname*{arg\,min}_{A}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{2}^{2}$ overcomes this correlation and achieves the consistency and the convergence towards the true matrix [ljung1976ols, moore1978strong, lai1982least]. Recently, instead of the asymptotic properties, several studies have focused on non-asymptotic analyses and provided an error rate of $O\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ between the true matrix and OLS estimates under i.i.d. zero-mean Gaussian and sub-Gaussian disturbances after finite samples are collected [simchowitz2018learningwithout, jedra2020finite, sarkar2021finite]. These results were extended to use the OLS estimates for the linear quadratic control with unknown matrices, in a way that the estimates are treated as the true matrices after a finite time, achieving $O(\sqrt{T})$ regret [mania2019certainty, simchowitz2020naive]. While OLS is indeed a statistically optimal estimator in such i.i.d. cases, little is known about the system identification when the disturbances are correlated or possibly selected adversarially based on the past information.
@@
Recently, in the machine learning literature, the online control problem with any bounded adversarial disturbance, called online nonstochastic control, has gained considerable attention, among which [hazan2020nonstochastic, simchowitz2020improper, chen2021blackbox] leveraged the OLS method to recover the matrix and attained $O(T^{2/3})$ regret. However, this regret implies a substantial cumulative loss, leading to an impractical algorithm for real-world applications. To address more applicable situations, [feng2021adv, yalcin2023exact, zhang2024exact] assumed that adversarial disturbances occur with probability less than 1, and recovered the true matrix with a non-smooth $l_{2}$-norm estimator defined by $\operatorname*{arg\,min}_{A}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{2}$. In this case, practicality is achieved since the regret does not grow after the exact recovery. However, they assumed that each nonzero adversarial disturbance has a zero mean, which overlooks the possibility of systemic bias in the disruptions.
@@
In this paper, we allow adversarial noncentral (nonzero-mean) disturbances and identify the unknown matrix within a finite time by using the $l_{1}$-norm estimator, given by $\operatorname*{arg\,min}_{A}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{1}$. Our work presents the first result in the literature for the exact retrieval of the true matrix even in the presence of adversarial noncentral disturbances. Under the assumptions that the probability of having an attack at a given time is $p\in(0,1)$ (see Assumption 3) and each attack covers the entire space in expectation (see Assumption 4), we construct two scenarios of the adversarial disturbance structure:
@@
    symmetric disturbance around zero scaled by a random vector (see Assumption 5) with $10<p<$ 1,
@@
    any adversarial noncentral disturbance with $0.50<p<$ 0.5.
The first scenario is applicable to the domain where positive and negative disturbances occur with equal chance, each of which can be scaled arbitrarily. For example, beliefs based on behavioral bias may cause fluctuations in the stock market [lo2004adaptive] or energy demand [werth2021energy], which can be explained by disturbances of arbitrary magnitude. A more practical situation arises in the second scenario: it applies when an extremely large attack with a nonzero mean occasionally affects the system, such as natural disaster on power systems [buldyrev2010cat, wang2016power, yan2020power], unanticipated malicious cyber attacks [pasqualetti2013cyber, duo2022cyber], and others. In this paper, we will show that the $l_{1}$-norm estimator is indeed a robust estimator capable of successfully prevailing over such disturbances. Note that while the paper focuses on autonomous linear systems, it can be readily generalized to linear systems with inputs as well as nonlinear systems linearly parameterized by basis functions. These types of generalization can be achieved by adopting the formulation and techniques studied in [yalcin2023exact, zhang2024exact], but this will be left as future work due to space restrictions.
@@
Notation. Let $\mathbb{R}^{d}$ denote the set of $d$-dimensional vectors and $\mathbb{R}^{d\times d}$ denote the set of $d\times d$ matrices. For a matrix $A$, $\|A\|_{2}$ denotes the operator norm and $\|A\|_{F}$ denotes the Frobenius norm of the matrix. For square matrices $A$ and $B$, $A\succeq B$ means that $A-B$ is positive semidefinite. For a vector $x$, $\|x\|_{1}$ denotes the $l_{1}$-norm of the vector and $\|x\|_{2}$ denotes the $l_{2}$-norm of the vector. $x^{T}$ denotes the transpose of the vector $x$. For two vectors $x$ and $y$ of the same dimensions, let $x\circ y$ denote entry-wise product of the same dimensions as $x$ and $y$. For a scalar $z\neq 0$, $\text{sgn}(z)=1$ if $z>0$ and $\text{sgn}(z)=-1$ if $z<0$. Let $\mathbb{E}$ denote the expectation operator. For the event $\mathcal{E}$, let $\mathbb{P}(\mathcal{E})$ denote the probability of the event. For the set $S$, let $|S|$ denote the cardinality of the set. We use $\Theta(\cdot)$ for the big-$\Theta$ notation, and $\tilde{\Theta}(\cdot)$ for the big-$\Theta$ notation hiding logarithmic factors. Finally, let $\mathbb{S}^{d-1}$ denote the set $\{y\in\mathbb{R}^{d}:\|y\|_{2}=1\}$.
@@ II Preliminaries
Definition 1 (sub-Gaussian scalar variables):
@@
Note that the $\psi_{2}$-norm satisfies all properties of norm, including positive definiteness, homogeneity, and triangle inequality. We have the following useful property of the moments of any sub-Gaussian variable $w\in\mathbb{R}$:
$\mathbb{E}[|w|^{p}]^{1/p}\leq 3\sqrt{p}\cdot\|w\|_{\psi_{2}},\quad\forall p\geq 1.$ (3)
We also introduce below an equivalent definition in the case where $\mathbb{E}[w]=0$.
Definition 2 (sub-Gaussian scalar variables (MGF)):
@@
$\mathbb{E}[\exp(\theta w)]\leq\exp(\theta^{2}\eta_{w}^{2})$ (4)
for all $\theta\in\mathbb{R}$. For a sub-Gaussian variable that satisfies (1) and $\mathbb{E}[w]=0$, (4) is satisfied with $\eta_{w}\leq 6\sqrt{e}\|w\|_{\psi_{2}}$ [vershynin2018high].
Real-world systems have a limit on their actuators so they cannot accept arbitrarily large inputs. This enables us to assume that each disturbance follows sub-Gaussian variables of which the tail event rarely occurs. The following lemma presents Hoeffding’s inequality [vershynin2018high], providing sharp bounds on the tail event for both Definitions 1 and 2.
Lemma 1 (Hoeffding’s inequality).
Suppose that a random variable $w$ satisfies (1) with its sub-Gaussian norm defined in (2). Then, for all $s>0$, we have
$\mathbb{P}(|w|\geq s)\leq 2\exp\Bigr{(}-\frac{s^{2}}{\|w\|_{\psi_{2}}^{2}}\Bigr{)}.$ (5)
In addition, if a random variable $w$ has zero mean and satisfies (4) with parameter $\eta_{w}$, then for all $s>0$, we have
$\mathbb{P}(w\geq s)\leq\exp\Bigr{(}-\frac{s^{2}}{(2\eta_{w})^{2}}\Bigr{)},$ (6a)
$\mathbb{P}(w\leq-s)\leq\exp\Bigr{(}-\frac{s^{2}}{(2\eta_{w})^{2}}\Bigr{)}.$ (6b)
@@
Definition 3 (sub-Gaussian vector variables):
@@
Now, the following lemma provides the connection between the $\psi_{2}$-norm and the operator norm.
Lemma 2.
For any matrix $A\in\mathbb{R}^{d_{1}\times d_{2}}$ and any sub-Gaussian vector $x\in\mathbb{R}^{d_{2}}$, we have $\|Ax\|_{\psi_{2}}\leq\|A\|_{2}\|x\|_{\psi_{2}}$.
Proof.
Given a matrix $A$ and a vector $x$, we have
$\|Ax\|_{\psi_{2}}$ $=\sup_{\|y\|_{2}\leq 1,y\in\mathbb{R}^{d_{1}}}\|x^{T}A^{T}y\|_{\psi_{2}}$
$\leq\sup_{\|A^{T}y\|_{2}\leq\|A\|_{2},y\in\mathbb{R}^{d_{1}}}\|x^{T}A^{T}y\|_{\psi_{2}}$
$\leq\sup_{\|z\|_{2}\leq\|A\|_{2},z\in\mathbb{R}^{d_{2}}}\|x^{T}z\|_{\psi_{2}}$
$=\sup_{\|\tilde{z}\|_{2}\leq 1,\tilde{z}\in\mathbb{R}^{d_{2}}}\Bigr{\|}\|A\|_{2}\cdot x^{T}\tilde{z}\Bigr{\|}_{\psi_{2}}=\|A\|_{2}\cdot\|x\|_{\psi_{2}},$
where the first inequality comes from the relationship
and the last equality is due to the homogeneity of the $\psi_{2}$-norm.
@@ III Problem Formulation
where $A^{*}\in\mathbb{R}^{d\times d}$ is the unknown true matrix, $x_{t}\in\mathbb{R}^{d}$ is the state, and $w_{t}\in\mathbb{R}^{d}$ is a disturbance injected into the system at time $t$. Our goal is to identify the true matrix $A^{*}$, given the state trajectory $x_{0},\dots,x_{T}$. We assume that $A^{*}$ has bounded operator norm and each $x_{0},w_{0},\dots,w_{T-1}$ is sub-Gaussian to prevent an unbounded growth of the system. We formally present the relevant assumptions below.
Assumption 1 (Operator norm):
@@
Assumption 2 (Maximum sub-Gaussian norm):
Define a filtration $\mathcal{F}_{t}=\bm{\sigma}\{x_{0},w_{0},\dots,w_{t-1}\}$. There exists $\sigma_{w}>0$ such that $\|x_{0}\|_{\psi_{2}}\leq\sigma_{w}$ and $\|w_{t}\|_{\psi_{2}}\leq\sigma_{w}$ conditioned on $\mathcal{F}_{t}$ for all $t\geq 0$ and $\mathcal{F}_{t}$.
Our major advancement over previous works is to address two main challenges simultaneously: the disturbances can be adversarial rather than independent, and each of them may also have a nonzero mean. This allows $w_{t}$ to be adversarial noncentral disturbances, meaning that an adversary can design the disturbance $w_{t}$ based on the previous information $\mathcal{F}_{t}$, and $\mathbb{E}[w_{t}~{}|~{}\mathcal{F}_{t}]$ is not necessarily zero.
Our goal is to construct an estimator that achieves a finite-time exact recovery, meaning that the error between the true matrix and the estimate should be exactly zero after a finite time. However, it is well-established that the lower bound of the error is $\Theta\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ if disturbances follow i.i.d. zero-mean Gaussian distribution [simchowitz2018learningwithout]. Evidently, this lower bound $\Theta\bigr{(}\frac{1}{\sqrt{T}}\bigr{)}$ does not allow exact recovery, and thus we introduce the following further assumption also adopted by [feng2021adv, yalcin2023exact, zhang2024exact].
Assumption 3 (Probabilistic Attack):
The disturbance has an attack probability $10<p<$ 1 conditioned on $\mathcal{F}_{t}$, meaning that
$\mathbb{P}(w_{t}\equiv 0~{}|~{}\mathcal{F}_{t})=1-p$ (9)
holds for all $t\geq 0$ and $\mathcal{F}_{t}$, where $\equiv$ means that the two sides are identically equal. We also define the attack time set as $\mathcal{K}_{T}=\{0\leq t\leq T-1:w_{t}\text{~{}is not identically zero}\}$.
Note that Assumption 3 implies that the system is not under attack at time $t$ with a positive probability $1-p$. We further require a similar assumption presented in [zhang2024exact]; When the system is under attack, no adversarial disturbance can deceive the system to the extent that the distribution of the next state fails to span the entire state space, thereby ensuring sufficient exploration of the state trajectory. Remark 2 provides more details on justifying such role of attacks.
Assumption 4 (Non-degeneracy):
There exists $\lambda>0$ such that for every $x\in\mathbb{R}^{d}$, it holds that
$\mathbb{E}[(x+w_{t})(x+w_{t})^{T}~{}|~{}\mathcal{F}_{t}]\succeq\lambda^{2}I_{d},$ (10)
for all $t\in\mathcal{K}_{T}$ and $\mathcal{F}_{t}$, where $I_{d}$ is the $d\times d$ identity matrix.
Now, given a state trajectory $x_{0},\dots,x_{T}$, we consider the following $l_{1}$-norm estimator at time $T$:
$\min_{A\in\mathbb{R}^{d\times d}}\sum_{t=0}^{T-1}\|x_{t+1}-Ax_{t}\|_{1}.$ (11)
We will show that $l_{1}$-estimator successfully overcomes adversarial noncentral distributions and achieves a finite-time exact recovery. We formally define this notion below.
Definition 4 (Finite-time Exact Recovery):
Let $S_{t}$ denote a set of solutions to the estimator based on $x_{0},x_{1},\dots,x_{t}$. The estimator is said to achieve finite-time exact recovery if for all $\delta\in(0,1]$, there exists $t_{\delta}>0$ such that
$t\geq t_{\delta}~{}\Rightarrow~{}S_{t}=\{A^{*}\}$ (12)
with probability at least $1-\delta$.
Note that Definition 4 implies that $A^{*}$ should be the unique solution to the estimator to achieve the exact recovery.